the perimeter around a rectangle is 48 inches. the length of the rectangel is 2 inches more thatn the width. what is the width of the rectangel?
Found 2 solutions by Manny Pacquiao, Jstrasner:
Answer by Manny Pacquiao(82)

2+2=4
22+22=44
+4=48
looks like 2 and 22 is your rectangle.. must be a weird looking rectangle..are you sure this is right?

Answer by Jstrasner(106)

Hey,
So for this one there are 2 equations that we have:
2L + 2w = 48
L = 2 + w
So all we need to do is substitute 2 + w for L in the first equation:
2(w+2) + 2w = 48 => 2w + 4 + 2w = 48 => 4w = 44 => w=11
Therefore:
L = 2 + 11 => 13
Therefore:
W= 11, L=13
I hope this helps!